AnimeGhoul8863:

I need some help with math i need someone to tutor me alittle bit so i can understand it better

Eiwoh2:

Well, nobody seems to be on right now that is an expert at math, so i'll tag @Elsa213 or @SmokeyBrown .

AnimeGhoul8863:

ok thx howie

Eiwoh2:

When they come on. And no problem.

563blackghost:

I can (depending what it is though)!

AnimeGhoul8863:

Algebra 2

563blackghost:

mm what subtopic?

AnimeGhoul8863:

ill put a question down that i dont know how to answer and we can work on it if thats ok

563blackghost:

ye sure

AnimeGhoul8863:

Given that f(x) = x2 + 2x + 3 and g(x) = quantity of x plus four, over three, solve for f(g(x)) when x = 2. 2 5 11 33 Im really bad with the whole f(g(x)) stuff

563blackghost:

When you are give f(g(x)) You will be plugging in your g(x) into the x-variable of f. So you have \(\bf{f(x)=x^{2} + 2x +3}\) For `every x` you plug in your equation of `g(x)`. Remember that we are told dat `x=2` so you would plug this in for `g(x)` \(\large\bf{g(x)=\frac{x+4}{3} \rightarrow \color{red}{\frac{2+4}{3}}}\) Simplify that and plug into each `x` of `f(x)`.

563blackghost:

still kinda lost?

AnimeGhoul8863:

yes how can x = g(x) so it would look like f(x) = g(x)2+2g(x) +3?

563blackghost:

ye it would! See you have `f(x)` and `g(x)` it tells you to simplify by `(f(g(x))` meaning for `x` of `f(x) it equals to g(x)`.

AnimeGhoul8863:

ok

AnimeGhoul8863:

so it looks confusing but truly its just the look of it

563blackghost:

ye cx there are these variables and complexities of it, but its simple to learn just gotta not look at it as a whole i guess, just in parts ;)

AnimeGhoul8863:

ok

AnimeGhoul8863:

i think i understand it alittle more